For a sample size of 100 people, what is the margin of error that 53% of consumers will vote “yes” on a new product line from Company A to Z?


1
Expert's answer
2021-05-24T15:07:41-0400

Given,

n = 100

Point estimate = sample proportion = "\hat p = 0.53"

"1-\hat p=1-0.53=0.47"

At 95% confidence level

"\alpha =1-95\%\\\alpha=1-0.95=0.05\\\alpha/2=0.025\\Z_{\alpha/2}=Z_{0.025}=1.96"


Margin of error = "E=Z_{\alpha/2}\sqrt{\dfrac{\hat p\cdot (1-\hat p)}{n}}"


"E= 1.96\sqrt{\dfrac{0.53\times 0.47}{100}}\\\Rightarrow E=1.96\times0.04991\\\Rightarrow E=0.0978"


"\boxed{\text{Margin of error = 0.0978}}"